OpenStudy (summersnow8):

K-1 whitewater kayak has a volume of 225 liters (0.225 m3) and a mass of 10 kg. Sean wonders if the kayak is big enough for him to use. Sean's mass is 70 kg. What percent of the kayak will be submerged when Sean sits in it? Assume that the water density is 1.00 g/cm3 (1000 kg/m3). Answer is 36% I need help solving

OpenStudy (anonymous):

i think your right

OpenStudy (summersnow8):

yeah, the answer is correct, but I dont know how to solve it, that is what I need help with @pi=3.14159265359

OpenStudy (anonymous):

ill look into it

OpenStudy (anonymous):

the kayak will stop sinking when the lift will be equal to the gravity. ill denote ms - mass of sean rk - density of kayak. Vk - volume of kayak. Vw - volume of displaced water which is in fact the volume of the kayak inside the water. rw- density of water so we are looking for Vw/Vk. ms * g + Vk * rk * g = Vw * rw * g Vk * rk is just mk (mass of the kayak) Vw = (ms + mk) / rw and Vw/Vk = (ms + mk / (rw*Vk) = 0.355 = 0.36 so the percentage (Vw/Vk) * 100% = 36%

OpenStudy (anonymous):

ms * g + Vk * rk * g = Vw * rw * g this is fine ?

OpenStudy (anonymous):

ms * g + Vk * rk * g = Vw * rw * g now i substitute Vk* rk = mk so ms * g +mk * g = Vw * rw * g we can divide g : ms +mk = Vw * rw solve for Vw : Vw = (ms + mk) / rw now divide both sides by Vk: Vw/Vk = (ms + mk) / (rw*Vk) plug numbers: Vw/Vk = (70 + 10)/ (1000*0.225) ...